OpenStudy (joftheworld):

What is the vertex of the parabola given as x2 − 4x − 2y = 0

OpenStudy (phi):

write \[ x^2 − 4x − 2y = 0 \\ 2y= x^2 -4x \\ y= \frac{1}{2}x^2 -2x \] the vertex of a parabola its equation in the form \[ y = ax^2 + bx + c\] is x= -b/(2a) use that value of x to find y using the equation y= ax^2 + bx + c